Solve the following systems by using either addition or substitution. If a unique solution does not exist, state whether the system is dependent or inconsistent.
38. 10x +2y = 7
Y=-5x +3

Answer by homeworkhelpanytime(21)

10x +2y = 7 ...(1)
Y=-5x +3
substitute the value of y in equation of (1)
10x + 2(-5x + 3) = 7
10x -10x +6 = 7
0 =1
The equations are incosistent